Graduation is 2 years away and you want to have $1450 available for a trip. If your bank is offering a 2-year CD (certificate of deposit) paying 2.4% simple interest, how much do you need to put in this CD to have the money for your trip?
Answer by addingup(3677)

P: Principal, the amount you need to put in the CD
r: interest rate
t: time
A: total amount
~~~~~~~~~~~~~~~~~~~~~~~
P(1+rt) = A
P(1+(0.024*2)) = 1450
P(1.048) = 1450
P = 1450/1.048 = 1383.59